OpenStudy (anonymous):

CHECK PROBLEM 3 PLEASE IM LOST (1) A baseball is thrown with a vertical velocity of 50 ft/s from an initial height of 6 ft. The height h in feet of the baseball can be modeled by h(t) = -16t2 + 50t + 6, where t is the time in seconds since the ball was thrown. It takes the ball approximately____seconds to reach its maximum height. (Round to the nearest tenth of a second and enter only the number.) I got 1.6 here. (2) To the nearest foot, what is the maximum height that the ball reaches? (Enter only the number.) I got 45 here.

OpenStudy (anonymous):

(3) A player hits a foul ball with an initial vertical velocity of 70 ft/s and an initial height of 5 feet. The maximum height reached by the ball is feet. (Round to the nearest foot and enter only the number.)

OpenStudy (jack1):

1.56 seconds so to the nearest tenth 1.6 is correct of Q1

OpenStudy (anonymous):

ok ty im pretty sure i got 1 and 2 right its the that has me lost

OpenStudy (jack1):

45.06 so round it to 45 ft is right for 2... now on to 3...

OpenStudy (jack1):

so what equation are u using to model height in this siutation...? becaust the other equation seems specific for q 1 and 2 only

OpenStudy (anonymous):

it didnt give me an equation i only got the question... maybe the equation from q1 is for q 3 idk...thats y im confused

OpenStudy (jack1):

no, that equation looks specific to a ball with a starting velocity of 50... Q3 's starting velocity is 70

OpenStudy (anonymous):

would i substitute the 50 for 70 in equation 1?

OpenStudy (anonymous):

-16t2 + 70t + 5 is what i think but idk

OpenStudy (jack1):

have u ever seen the equation s = ut + 1/2 at^2 before...? \[s = ut + \frac{ 1 }{ 2 }at^2\]

OpenStudy (anonymous):

no i havent

OpenStudy (jack1):

ok, well it seems to fit, as per your above post a = gravity = -32.152 ft/second u = 70 so s = ut + 1/2 at^2 ht = 70t + 1/2 * 32 * t^2 ht = 70t - 16t^2 .... +5ft of initial height

OpenStudy (jack1):

so yeah, use this equation: "-16t2 + 70t + 5 is what i think but idk"

OpenStudy (anonymous):

ok thanks ill try it

OpenStudy (jack1):

post ur answers if u could dude, we'll compare...?

OpenStudy (anonymous):

i got 5? im not sure if i typed it in right tho

OpenStudy (jack1):

for time or height...? because i got less than half that for time, and whay more for height

OpenStudy (anonymous):

nvm i got 311.25 for height 4.375 for time

OpenStudy (jack1):

...whoa... that's huge dude, some bloke smashed a ball 300m straight up... ;) so what eqn jew use work out time?

OpenStudy (anonymous):

h(t) = -16t^2 + 50t +6 i did h/t= 70/16 and got 4.375 for t and just plugged that in t

OpenStudy (jack1):

ok, it should be 70t tho... ht = -16t2 + 70t + 5 and h/t... i dont really understand that dude, sorry I did it the same way i did Q1... of i got the equation: h(t) = -16t2 + 70t + 5 and i found the derivative h'(t) = -32t + 70 so the derivative is the gradient at any point in time so when the gradient = 0, its a turning point (as your initial function is a parabola) so h'(t) = -32t + 70 0 = -32t + 70 32t = 70 so t = 70/32 = 2.1875 seconds

OpenStudy (jack1):